Fox Hollow Franks is looking at a new system with an installed cost of $540,000. This equipment is depreciated at a rate of 20% per year (Class 8) over the project’s five-year life, at the end of which the sausage system can be sold for $80,000. The sausage system will save the firm $170,000 per year in pre-tax operating costs, and the system requires an initial investment in net working capital of $29,000. If the tax rate is 34% and the discount rate is 10%, what is the NPV of this project? (Do not round your intermediate calculations. Round the final answer to 2 decimal places. Omit $ sign in your response.)

 

NPV 
Answer #1

Please see the table below. Please be guided by the second column titled “Linkage” to understand the mathematics. The last row highlighted in yellow is your answer. Figures in parenthesis, if any, mean negative values. All financials are in $. 

Year, nLinkage012345
Asset costA540000




Class 8 depreciation rated20%




Opening balance

   540,000        432,000         345,600    276,480       221,184
DepreciationD
   108,000          86,400           69,120      55,296          44,237
Closing balance

   432,000        345,600         276,480    221,184       176,947








Sale ValueS




          80,000
Gain / (Loss) on saleB = S - closing balance of year 5




        (96,947)
TaxT = B x 34%




        (32,962)
Post tax salvage valueC = S - T




       112,962








Cash flows






Asset costA   (540,000)




Savings in pre tax operating costE
   170,000        170,000         170,000    170,000       170,000
[-] DepreciationD
   108,000          86,400           69,120      55,296          44,237
EBITF = E - D
     62,000          83,600         100,880    114,704       125,763
NOPATG = F x (1 - 34%)
     40,920          55,176           66,581      75,705          83,004
OCFH = G + D
   148,920        141,576         135,701    131,001       127,241
Working capitalI     (29,000)



          29,000
Post tax salvage valueC from above




       112,962








Net cash flowsCF = A + H + I + C   (569,000)   148,920        141,576         135,701    131,001       269,203
Discount rater10%




PV factorPVF = (1 + r)^(-n)1.00000.90910.82640.75130.68300.6209
PV of cash flowsCF x PVF   (569,000)   135,382        117,005         101,954      89,475       167,154
NPVSum of all PVs 41,969.61





answered by: Book Solutions